If a vendor charges p dollars each for rugby shirts, then he expects to sell 2000-100p shirts at a tournament.
a) Find the polynomial R(p) that represents the total revenue when the shirts are p dollars each.
b)Find R(5), R(10), and R(20)

Answer by stanbon(75887)

If a vendor charges p dollars each for rugby shirts, then he expects to sell 2000-100p shirts at a tournament.
a) Find the polynomial R(p) that represents the total revenue when the shirts are p dollars each.
R(p) = p(2000-100p)
----------------------------
b)Find
R(5) = 5(2000-500) = 7500
R(10)= 10(2000-1000) = 10000
R(20)= 20(2000-2000) = 0
==============================
Cheers,
Stan H.